INAF LBC+LUCIFER queue observing, 2011 June 26 -27

Observer: R. Speziali, K. Boutsia
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LBC

Summary

Sky is partly clouded at the beginning of the night. We decide to go on with LBC anyway, because according to weather forecast it might get worse in the following days.
Trouble at the beginning of the night because the pupils were not the right size for dofpia to work properly. At some point the blue camera turned out to be intrafocal. It took 1h45min to collimate with dofpia. These problems are probably related to the temperature being too warm for the collimation model to handle correctly and the blue mirror being >1deg colder than the ambient temperature.

Data Summary

Project Objects Filter ExpTime Notes
Lanzoni Pal14_F1 g-SLOAN - r-SLOAN 1hr it should be within specs - some images a bit elongated - had to stop because of clouds
Magrini 7c17_cluster 210_zJHK - J 2hr15min arc+flat + Tell (?)

Details

02:14 - Bias frames with the LBC not connected to the telescope while the T.O switches configuration
  • 20Bias_Bino, 5Bias_Bino_rot90, 5Bias_Bino_rot180, 5Bias_Bino_rot270

02:54 - Slewing for twilight flats in g-SLOAN and r-SLOAN
  • the sky is partly covered with clouds and we don't know if flats will be good
  • in case tonight's flats are not usable, flats in g-SLOAN and r-SLOAN have been obtained during the SciPrep night (UT - 20110624)

03:16 - Slew to ACT0247 for collimation and pointing correction
  • add Z11=+1000 at the blue side
  • dofpia, /First - at 2nd iteration , no good pupils found
  • dofpia, /Backout
  • dofpia, /First, /X2
  • red is converged but blue fails - dofpia exited after 8 iterations
  • Z4=-20000 in the blue and we repeat dofpia, /Blueonly - exited after 7 iterations
  • dofpia - again - it exited after the 3rd iteration because red is converged and cannot find good pupils in the blue channel

04:14 - lbcrangebal to correct pointing

04:17 - dofpia again but this time we got an RPC Client/Host communication error
  • Red camera encountered an error
  • Turn OFF/ON LBC
  • the mirror is about 1 colder than the ambient

04:28 - camera restarted - Slew to ACT0247- dofpia
  • dofpia is failing
  • Olga is taking a look at the images
  • called Andrew at 04:10 (???)

04:36 - dofpia again after reconnecting LBC to the telescope
  • before it seemed that we were intrafocus and that was a problem
  • we gave some negative Z4 - now we have a big amount of coma
  • we correct coma Z7= -2000 - the spider is on the same side in both images and we run dofpia

04:50 - dofpia * it seems to be working now... andrew out

04:58 - Slew to ACT0247 for rangebal

05:02 - slew to Pal14_F1_1
  • dofpia

05:07 - Pal14_F1_1
  • seeing measured on first image is ~1.1'' for the blue and <1.0'' for the red
  • DIMM seeing 1.5''

05:40 - dofpia
  • it took 2 iterations to converge

05:44 - Pal14_F1_2
  • images seem somewhat elongated
  • seeing measured on these images ~1.3''

06:17 - dofpia
  • failed because it cannot find pupils
  • clouds are getting thicker
  • dofpia, /X2 - one more try...
  • we can see no pupils in the dofpia - there are clouds

06:23 - Switch to Lucifer

06:50 - pointing correction

07:00 - collimation
  • it is very difficult to collimate because the star is barely visible - probably because of the clouds
  • choosing a brighter star for collimation

07:44 - magrini_tell_acq 0003-000?
  • we have changed the guide star with a brighter one

07:50 - magrini_tell_int_J ???

08:00 - magrini_acq_sci.txt 0012-0015
  • doing MOS acquisition

08:20 - while we were sending the offsets obtained during MOS acq the preset failed
  • guide star was lost - probably due to the clouds

08:21 - magrini_acq_sci.txt - error committing the telescope

08:22 - correct pointing

08:26 - magrini_acq_sci.txt 0016-0021

08:49 - magrini_spec_sci_J_2.txt 0022-0033
  • DIMM seeing 0.9''- guider also

09:56 - magrini_spec_sci_J_3.txt 0034-0045

11:02 - magrini_spec_sci_J_4.txt 0046-0049
  • after the 3rd offset we interrupted the script because it is getting bright

11:20 - Calib Unit IN
  • ARC 0050-0060
  • FLAT 0061-0075

11:30 - magrini_tell_acq - preset failed
  • check pointing
  • calib unit was IN

11:36 Calib Unit OUT

11:40 - magrini_tell_acq
  • the guide star is in the GCS for an instant and then it slews away
  • repeat acquisition and turn off guiding

11:46 - magrini_tell_int_J.txt 0081-0084
  • we have changed the offsets in the script in order to put the telluric in a slit
  • it seems that the telluric in the second slit is saturated (why?? - in the first it has reasonable counts...)
  • we reverse the offsets and obtained the tell again

11:51 - magrini_tell_int_J.txt 0085-0088
  • this time it seems that the telluric is not well centered - maybe because we are not guiding
  • we try to center manually and take an image outside the script 0089
  • still very low counts ... probably telluric is drifting away
  • we stop because it is getting bright

11:55 - End of night - DoDarks

-- KonstantinaBoutsia - 27 Jun 2011
Topic revision: r5 - 27 Jun 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback